### v2

# Scrambled Halton Subsequences and Inverse Star-Discrepancy

Abstract
Braaten and Weller discovered that the star-discrepancy of Halton sequences can be strongly reduced by scrambling them. In this paper, we apply a similar approach to those subsequences of Halton sequences which can be identified to have low-discrepancy by results from $p$-adic discrepancy theory. For given finite $N$, it turns out that the star-discrepancy of these sequences is surprisingly low. By that known empiric bounds for the inverse star-discrepancy can be improved. Furthermore, we establish the existence of $N$-point sets in dimension $d$ whose star-discrepancy satisfies $\leq 2.4631832\sqrt{\frac{d}{N}}$, where the constant improves upon all previously known bounds.
1 Introduction
Let $(x_{n})\subset[0,1]$ be a uniformly distributed sequence, i.e. for all $10\leq a<b<$ 1 it holds
$\lim_{N\to\infty}\frac{\#\left\{1\leq n\leq N\,:\,x_{n}\in[a,b)\right\}}{N}=b-a.$
It is well-known that for every uniformly distributed sequence $(x_{n})$ there exists a re-ordering of its elements $(y_{n}):=(x_{\sigma(n)})$ by some bijection $\sigma:\mathbb{N}\to\mathbb{N}$ such that its star-discrepancy
$D_{N}^{*}(y_{n}):=\sup_{0<b\leq 1}\left|\frac{\#\left\{1\leq n\leq N\,:\,y_{n}\in[0,b)\right\}}{N}-b\right|$
is of the best possible order $D_{N}^{*}(y_{n})=O(\log(N)/N)$, see [Sch72]. Sequences satisfying this asymptotic property are called low-discrepancy sequences. Similarly, the (extreme) discrepancy is defined by
$D_{N}(y_{n}):=\sup_{0\leq a<b\leq 1}\left|\frac{\#\left\{1\leq n\leq N\,:\,y_{n}\in[a,b)\right\}}{N}-(b-a)\right|,$
i.e. the intervals in the supremum are not necessarily anchored at 00. It relates to the star-discrepancy via $D_{N}^{*}(y_{n})\leq D_{N}(y_{n})\leq 2D_{N}^{*}(y_{n})$. Thus, the two notions necessarily possess the same asymptotic behavior up to a factor.
[…]
Recall that for an integer $b\geq 2$ the $b$-ary representation of $n\in{\mathbb{N}}$ is $n=\sum_{j=0}^{\infty}e_{j}b^{j}$ with $0\leq e_{j}=e_{j}(n)<b$. Based on the radical-inverse function, the van der Corput sequence in base $b$ is defined by $\varphi_{b}(n)=\sum_{j=0}^{\infty}e_{j}b^{-j-1}$ for all $n\in{\mathbb{N}}_{0}$. It is well-known, see e.g. [Nie92], that
$D_{N}^{*}(\varphi_{b}(n))\leq\frac{1}{N}+\frac{b+1}{2N}+\frac{b-1}{2\log(b)}\frac{\log(N)}{N}$
This bound should be compared to (and is for small $b$ not too far off from) the record holder for the best known star-discrepancy as constructed in [Ost09] which satisfies
$\limsup_{N\to\infty}\frac{ND_{N}^{*}(x_{n})}{\log(N)}\approx 0.222223.$
In fact, this record holder is a generalized van der Corput sequence in base $b=60$. In other words, van der Corput, their subsequences and generalizations may be regarded as prime candidates when looking for sequences with a particularly small star-discrepancy.
An easier-to-describe method than in [Ost09] how to (empirically) further reduce the star-discrepancy of van der Corput sequences was analyzed in detail in [BW79] and was based on an idea from [Fau78]: for fixed $b\in\mathbb{N}$ choose an arbitrary permutation $\pi$ of $\{0,\ldots,b-1\}$ and define the scrambled van der Corput sequence by
$\varphi_{b,\pi}(n):=\sum_{j=0}^{\infty}\pi(e_{j})b^{-j-1}.$ (1)
It is not difficult to prove the same bound for the star-discrepancy as for the standard van der Corput sequence but the additional parameter allows for (empiric) reduction of the star-discrepancy. In [BW79], concrete choices for $\pi$ were suggested for all prime bases $p\leq 53$.
Actually, the result in [Ost09] relies on the (rather complicated) exact formulae for the star-discrepancy of scrambled van der Corput sequences derived in [Fau81]. The article [Fau81] also provides asymptotic constants for the limit superior of both the extreme discrepancy and the star discrepancy. As another main theoretical result it was proven in [Fau92], that for every base $b$ there exists a permutation $\pi$ such that the constant in the inequality $D_{N}^{*}(\varphi_{b,\pi}(n))\leq c\frac{\log(N)}{N}$ is less than $1/\log(2)$.
The concept of scrambling can, of course, also be applied to subsequences of van der Corput sequences and we obtain the following result which may be regarded as a generalization of Theorem 1.2 in [Wei25] in combination with Theorem 3 in [Mei68] as will become clear from its proof. Recall that a polynomial $f:\mathbb{Z}\to\mathbb{Z}$ is called a permutation polynomial $\textrm{mod}\ n$ for $n\in\mathbb{N}$ if it is a bijection on $\mathbb{Z}/n\mathbb{Z}$ and thus a permutation of the elements in $\mathbb{Z}/n\mathbb{Z}$.
Theorem 1.
Let $f(n)$ be a permutation polynomial $\bmod\ p^{2}$ for some prime number $p$, which means that $f(n)$ induces a bijection on $\mathbb{Z}/p^{2}\mathbb{Z}$, and let $\pi$ be a permutation of $\{0,\ldots,p-1\}$. Then the extreme discrepancy of the sequence $(x_{n}):=(\varphi_{p,\pi}(f(n)))$ satisfies
$D_{N}(x_{n})\leq\frac{p-1}{2\log(p)}\frac{\log(N)}{N}+O\left(\frac{1}{N}\right).$
We call such sequences scrambled van der Corput subsequences. The easiest examples are of the form $f(n)=an+b$ with $\gcd(a,p)=1$. We call $a$ a shift and denote such a sequence by $\varphi_{p,\pi,a}$. Other possible choices for $f(n)$, depending on $p$, may be found in [Wei25]. Although the theoretical bound does not guarantee a very small extreme or star-discrepancy, allowing the flexibility in choosing the parameters $\pi$ and $a$ in Theorem 1 reduces the extreme or star-discrepancy for given $N$ significantly as can be seen from Table 1 where our results are compared to the approach from [BW79]. The difference between the original van der Corput sequence and the scrambled one according to [BW79] is much larger (except for $p=2$), but moving to subsequences always but once reduced the star-discrepancy as well. In our experiments, the gain exceeded $10\%$ most of the time.
p $\varphi_{p}(N)$ $\varphi_{p,\pi}(N)$ $\varphi_{p,\pi,a}(N)$
2 0.0231 0.0231 0.0143
3 0.0262 0.0199 0.0140
5 0.0160 0.0120 0.0100
7 0.0310 0.0182 0.0107
11 0.0321 0.0198 0.0111
13 0.0563 0.0189 0.0137
17 0.0503 0.0144 0.0121
19 0.0731 0.0153 0.0125
23 0.0846 0.0150 0.0144
29 0.0982 0.0167 0.0162
p $\varphi_{p}(N)$ $\varphi_{p,\pi}(N)$ $\varphi_{p,\pi,a}(N)$
2 0.0025 0.0025 0,0018
3 0.0031 0.0028 0,0018
5 0.0025 0.0018 0,0010
7 0.0042 0.0023 0,0018
11 0.0049 0.0024 0,0018
13 0.0046 0.0019 0,0017
17 0.0086 0.0022 0,0021
19 0.0095 0.0019 0,0020
23 0.0093 0.0022 0,0020
29 0.0123 0.0025 0,0023
Table 1: Comparison of star-discrepancies for van der Corput like sequences for $N=100$ (left) and $N=1,000$ (right). For $\varphi_{p,\pi,a}(N)$ the smallest found values are listed when using $500$ different shifts, $500$ random permutations per shift.
Finding sequences with a particularly small (star-)discrepancy is not only a relevant question in dimension $d=1$ but even of higher importance for $d>1$, where the star-discrepancy for a sequence $(x_{n})\subset[0,1]^{d}$ and the $d$-dimensional Lebesgue measure $\lambda_{d}(\cdot)$ is defined by
$D_{N}^{*}(x_{n}):=\sup_{B\subset[0,1)^{d}}\left|\frac{\#\left\{1\leq n\leq N\,:\,x_{n}\in B\right\}}{N}-\lambda_{d}(B)\right|,$
where the supremum is taken over all $d$-dimensional intervals $B=[0,b_{1})\times\ldots\times[0,b_{d})$ with $0\leq b_{i}\leq 1$ for $i=1,\ldots,d$. It is widely conjectured that $D_{N}^{*}(x_{n})=O(\log(N)^{d}/N)$ is the best achievable order of convergence for the star-discrepancy but this conjecture has only been proven in the case $d=1$, see [Sch72]. Furthermore, we remind the reader that the star-discrepancy of sequences (infinitely many points) in dimension $d$ corresponds to that of point sets (finitely many points) in dimension $d+1$, see [Rot54]. The extreme discrepancy $D_{N}(x_{n})$ again allows arbitrary multi-dimensional half-open intervals in the supremum, i.e. without necessarily anchoring one base point at zero. As a generalization of the one-dimensional case, the inequality $D_{N}^{*}(x_{n})\leq D_{N}(x_{n})\leq 2^{d}D_{N}^{*}(x_{n})$ holds.
Sequences with a particularly small (star-)discrepancy are of great interest for high-dimensional integration tasks. Due to the Koksma-Hlawka inequality, see e.g. [KN74], the worst case error when approximating an integral by the average of some function evaluations depends linearly on the star-discrepancy of the evaluation points. This motivates why it makes sense to ask, what is the smallest star-discrepancy achievable for a given $N\in\mathbb{N}$ and to set
$D^{*}(N,d):=\inf\{D_{N}^{*}(P)\,:\,P\subset[0,1]^{d},\#P=N\}.$
Furthermore, define the inverse star-discrepancy by
$N^{*}(\varepsilon,d):=\inf\{N\in\mathbb{N}\,:\,D^{*}(N,d)\leq\varepsilon\},$
which is the minimum number of sample points that guarantees a star-discrepancy bound of at most $\varepsilon>0$. Note that even small reductions of $N$ are of practical importance if an expensive function $f$ is evaluated at these points to (numerically) calculate an integral. Alternatively, decreasing the size of $N$ can be motivated by the numerical stability of certain regression problems which depend linearly on the star-discrepancy, see [WN19]. Therefore, precise theoretical bounds and numerical estimates of $N^{*}(\varepsilon,d)$ are of great interest.
As the asymptotic bound $\log(N)^{d}/N$ for the star-discrepancy exponentially depends on the dimension, the enumerator is large in comparison to the denominator if $N$ is small. Therefore, the bound is hardly of any use in this setting. In the seminal paper, [HNWW00], an alternative upper bound for the smallest achievable star-discrepancy of the form
$D^{*}(N,d)\leq c\sqrt{\frac{d}{N}}$ (2)
for some constant $c>0$ was shown without giving an explicit value for $c$. This implies
$N^{*}(\varepsilon,d)\leq\lfloor c^{2}d\varepsilon^{-2}\rfloor$
for all $d,N\in\mathbb{N}$ and $\varepsilon\in[0,1)$. Currently, the best known value for the constant is $c=2.4968$ according to [GPW21]. This observation can also be used to check whether a given point set or sequence is good in the sense that its empirically observed star-discrepancy is close to or even smaller than the upper bound. As a special example for this idea, it was shown in [GGPP21] that the sequence generated by a secure bit generator is good up to dimension at least $d=15$, because its star-discrepancy is smaller than $\sqrt{\frac{d}{N}}$ even for relatively large $N$.
Finding Multi-Dimensional Sets with Small Star-Discrepancy.
A classical example of multi-dimensional low-discrepancy sequences are the so-called Halton sequences which are the main object of study in this paper: for a given dimension $d$, let $b_{1},\ldots,b_{d}$ be pairwise relatively prime integers. The Halton sequence $(H^{b}_{n})$ in the base $b=(b_{1},\ldots,b_{d})$ is given by $x_{n}:=(\varphi_{b_{1}}(n),\ldots,\varphi_{b_{d}}(n))$ for all $n\geq 1$. Scrambled Halton sequences are then defined by choosing a permutation $\pi_{b_{i}}$ for each $i\in\{1,\ldots,d\}$. Theorem 1 easily generalizes from van der Corput sequences to higher dimensions by the work of Meijer. Indeed, Theorem 5 in [Mei68] can be applied to obtain the following version in several dimensions.
Theorem 2.
Let $p_{1},\ldots,p_{d}$ be distinct prime numbers and let $\pi_{i}$ be an arbitrary permutation of $\{0,\ldots,p_{i}-1\}$ for each $i=1,\ldots,d$. Moreover, let $f_{i}(n):\mathbb{Z}\to\mathbb{Z}$ be a permutation polynomial $\bmod\ p_{i}^{2}$. Then
$(x_{n}):=(\varphi_{p_{1},\pi_{1}}(f_{1}(n)),\ldots,\varphi_{p_{d},\pi_{d}}(f_{d}(n)))$
satisfies
$D_{N}(x)\leq\frac{\log(N)^{d}}{N}\prod_{j=1}^{d}\frac{2(p_{i}-1)}{\log(p_{i})}+O\left(\frac{\log(N)^{d-1}}{N}\right).$
Thus, all scrambled Halton subsequences of the form $f(n)=(a_{1}\cdot n,a_{2}\cdot n,\ldots a_{d}\cdot n)$ with $\gcd(a_{i},p_{i})=1$ for $i=1,\ldots,d$ are low-discrepancy sequences. We also call these $a_{i}$ shifts.
Remark 3.
Note that for the original Halton sequence, the pre-factor in front of $\log(N)^{d}/N$ is known to be smaller by $2^{2d}$, if we look at the star-discrepancy instead of the extreme discrepancy, see e.g. [KN74].
Also in the multi-dimensional setting, the permutations $\pi_{i}$ and the shifts $a_{i}$ may be chosen in order to minimize the star-discrepancy. Our numerical results in this setting are promising and we obtain values for the star-discrepancy which are at least comparable to other recent methods for finding sets with small star-discrepancy, in particular[DR13], [CVN+23]. As we want to analyze this case extensively but not overload the introduction, we postpone the detailed discussion to Section 2.
A mathematical application.
Since finding (empiric) small values for the (inverse) star-discrepancy requires an extensive search, it can only cover some cases $N,d\in\mathbb{N}$. As a theoretical application of scrambled Halton (sub-)sequences, they can be used to improve the value of $c$ in the theoretical bound (2). In fact, numerical calculations for finitely many $N$ can cover those cases for which an application of bounds like in Theorem 2 is not sufficient.
Theorem 4.
For any $d,N\in\mathbb{N}$ we have
$D^{*}(N,d)\leq 2.4631832\sqrt{\frac{d}{N}}.$
The general structure of the proof for Theorem 4 is in parallel to [GPW21]. However, we add some additional ideas to improve the bound: First, we were able to slightly sharpen the arguments to derive the bounds. Second, we include a very recent result on bracketing numbers from [Gne24]. Last but not least, we employ Halton sequences along with improved bounds on their star-discrepancy from [Ata04] to address the case $d\leq 4$ whereas previously only the case $d=1$ had been treated separately. It turns out that the star-discrepancy of finitely many Halton sequence points must be computed individually for each dimension $d$ to complete the proof. Without utilizing the result from [Ata04] the number of necessary computations in dimension 4 would be prohibitively large. Even with this result, additional arguments are needed to further reduce the computational effort, making it feasible to carry out the calculations on a standard computer. This double reduction in computational complexity constitutes the main new contribution of the present article. We will discuss the details in Section 3, see in particular Remark 13.
It is natural to ask why our approach is only applied to the case $d\leq d_{0}=4$. From a theoretic viewpoint, there is no reason for this choice and we would expect that (scrambled) Halton (sub-)sequences could be used for any finite $d_{0}$: given $d_{1}\leq d_{0}$, the bound from [Ata04] can be applied to prove an analogue of Theorem 4 for all but finitely many $N\in\mathbb{N}$. However, these finitely many exceptions need to be checked on a computer, which is a very demanding task, especially in higher dimensions.
In [GSW09] it was proven that calculating the (star-)discrepancy is an NP-hard problem. Indeed, all known algorithms for calculating the star-discrepancy have exponential run time. The currently best known one was introduced in [DEM96] and is also called DEM algorithm named after its inventors Dobkin, Epstein and Mitchell. Its runtime is of the order $O(N^{1+d/2})$, where $N$ is the cardinality of the set, compare [CVN+23]. For our calculations in higher dimensions, see Section 2, we use a more recent implementation of the DEM algorithm based on earlier work of Magnus Wahlström publicly available under [CVdN+23]. We do not describe its details here but refer to the description of the DEM algorithm in [DGW14].
As the runtime of the DEM algorithm is of order $O(N^{d/2+1})$ our approach seems to be infeasible from some dimension on. If we assume that it is conducted up to e.g. $d_{0}=9$, then the constant $c$ would go down to approximately $2.4543$.
For higher dimensions and relatively large $N$, also the (exact) DEM algorithm becomes infeasible and the star-discrepancy can only be estimated approximately then. In this case, the Threshold Accepting (TA) algorithm is a good solution. It was originally described in [WF97] and later on improved in [GWW12], see also [DGW14]. An implementation of the TA algorithm used to check the calculations of this article is provided under the same link as the DEM algorithm, [CVdN+23].
The remainder of the paper is organized as follows: In Section 2, we discuss numerical results in a multi-dimensional setting, which are based on Theorem 2. We will show for many different combinations of $d,N\in\mathbb{N}$ that we obtain sets with a star-discrepancy which is at most as big as the currently best known ones obtained by alternative methods. Afterwards, we will give proofs of our theoretical results in Section 3. The proofs of Theorems 1 and Theorem 2 rely on $p$-adic discrepancy theory as discussed in e.g. [Wei25]. Furthermore, Theorem 4 is obtained by using Halton (sub-)sequences in small dimensions and an improved bound on bracketing numbers proven in [Gne24]. Finally, we collect the explicit combinations of primes, shifts and permutations by which we obtain our optimal numerical results in the Appendix 4. This puts the reader into the position to reproduce our calculations.
Acknowledgments.
The author would like to thank Francois Clément for discussions on the content of the paper and for providing the link to the C-code used to perform the calculations of the DEM algorithm. Moreover, he is grateful to Michael Gnewuch for his comments on a preliminary version of this paper and especially on the proof of Theorem 4.
2 Some Practical Applications
Theorem 2 yields a class of low-discrepancy sequences with three different parameters which can be chosen to minimize the star-discrepancy. These are the primes, the permutations and the shifts of the scrambled Halton subsequences. They allow to empirically minimize the star-discrepancy for a given $N$ or equivalently minimize the inverse star-discrepancy. We will see that even much smaller values than $\frac{\log(N)^{d}}{N}\prod\frac{p_{i}-1}{2\log(p_{i})}$, which would be the best we can hope for according to Theorem 2 when ignoring the $O(\tfrac{\log(N)^{d-1}}{N})$ term, can be realized this way.
One-dimensional sequences.
At first, we again discuss the one-dimensional situation. For a sound comparison of our approach with the original and the scrambled van der Corput sequence it should be born in mind that the star-discrepancy is by our method minimized for a given $N$ which might mean that it is particularly small for the chosen $N$ but not for many other $n<N$. In Figure 1, we therefore plot the star-discrepancy values of the usual van der Corput sequence $\varphi_{3}$, the scrambled van der Corput sequence $\varphi_{3,\pi}$ according to [BW79] and the best scrambled van der Corput subsequence $\varphi_{3,\pi,a}$ for $N=1,000$ according to our approach for $0001\leq n\leq 1,$ 000. The values for the star-discrepancy are rescaled by $n/\log(n)$ so that the plots remain readable throughout and the asymptotic behavior is better visible. In Figure 2, the same is done for $p=5$ instead. The scrambled sequence as well as the scrambled subsequence almost systematically outperform the original van der Corput sequence, while the former two methods have the lowest star-discrepancy for approximately the same number of $n\leq 1,000$.
Figure 1: Comparison of $\varphi_{3},\varphi_{3,\pi}$ and $\varphi_{3,\pi,1,000}$ up to $N=1,000$.
Figure 2: Comparison of $\varphi_{5},\varphi_{5,\pi}$ and $\varphi_{5,\pi,1,000}$ up to $N=1,000$.
An additional interesting question concerns how well the sequences $\varphi_{p,\pi,a}$ perform beyond the anchoring value $N$. In Figure 4 of the appendix, we expanded Figure 2 up to $N=5,000$ and the empirical observations remain the same. The sequences $\varphi_{\pi,5}$ and $\varphi_{\pi,5,a}$ have the lowest star-discrepancy for about the same number of $n$ and both clearly outperform the original van der Corput sequence. Similar observations can be made for other values of $p$.
In a more comprehensive setting, the star-discrepancy of scrambled van der Corput sequences may also be compared to other low-discrepancy sequences. We would like to address here two specific examples which both are known to have a particularly good asymptotic behavior, see e.g. [Clé24]. The first one is the golden ratio Kronecker sequence defined by $(x_{n})=\{n\phi\}$, where $\phi=\frac{1+\sqrt{5}}{2}$ is the golden ratio. The other, more recent one, is the Kritzinger-sequence $Kri_{n}$ introduced in [Kri21]: set $Kri_{1}=\frac{1}{2}$ and define $Kri_{n+1}$ for $n\geq 1$ by
$Kri_{n+1}=\arg\min_{x\in[0,1)]}-2\sum_{k=1}^{n}\max(Kri_{k},x)+(x+1)x^{2}-x.$
If there are several solutions, then always the smallest one is chosen as $Kri_{n+1}$. It can then be shown that all elements are of the form $Kri_{n}=\frac{2k-1}{2n}$ for some $1\leq k\leq n$.
In Figure 3, we compare the rescaled star-discrepancy of these two sequences with $\varphi_{5,\pi,a}(1000)$ and $\varphi_{5,\pi,a}(2000)$ up to $n=2,000$, where the number in the bracket indicates the anchoring value $N$ of the scrambled van der Corput subsequences. It turns out that $\varphi_{5,\pi,a}(1000)$ has the lowest star-discrepancy for $150$ different values of $n$ and $\varphi_{5,\pi,a}(2000)$ for $475$ different values while the best value is achieved for the golden ratio Kronecker sequence in $637$ cases and for the Kritzinger sequence in $737$. In a direct comparison of $\varphi_{5,\pi,a}(2000)$ and the Kritzinger sequence, $Kri_{n}$ wins $1,250$ times. However, it is not surprising that $\varphi_{5,\pi,a}(2000)$ is the winner of $80$ of the last $100$ values of $n$. Similarly, $\varphi_{5,\pi,a}(1000)$ is the winner of all four sequences $41$ times in the range $n=901,\ldots,1.100$. Summing up, scrambled van der Corput sequences seem to be good candidates when searching for sequences with a particularly small local star-discrepancy in a given (small) range but they also yield good star-discrepancy values for other $n$.
Figure 3: Comparison of $\varphi_{5},\varphi_{5,\pi}$ and $\varphi_{5,\pi,1,000}$ up to $N=2,000$.
Multi-dimensional setting.
To further prove the use of our method, we compare more numerical results to related work in the multi-dimensional case. To the best of the authors’ knowledge the best empiric results on the inverse star-discrepancy were very recently achieved in [CDP23] by the so-called subset selection algorithm. It provides an heuristic algorithm to find the $m$-element subset $B$ of a larger set $A\subset[0,1]^{d}$ with $|A|=n$ such that the star-discrepancy is minimal. The algorithm is swap-based and attempts to replace a point of the current subset by one currently not chosen, using the box with the worst local discrepancy.
Two other particularly relevant approaches to empirically minimize the inverse star-discrepancy are the following: while evolutionary algorithms are applied in [DGW10], a machine learning algorithm is used to generate a new class of low-discrepancy point sets named Message-Passing Monte Carlo (MPMC) points in [RKB+24].
We compare our results to the values in [CDP23], Table 1, first. We look at dimensions $4$ and $5$ and search for scrambled Halton subsequences which minimize the inverse star-discrepancy. We see in Table 2, that our algorithm yields smaller values for the inverse star-discrepancy than in [CDP23]. To be fair, we stress, however, that the results from the subset selection algorithm were not explicitly optimized to get particularly small values for the inverse star-discrepancy for all small $N\in\mathbb{N}$ but only some specific values $N$. So there might be some further scope for improvement using this method.
Dimension Target $D_{N}$ Sobol’ Subset Selection Scrambled Halton
d = 4 0.30 15 10 8
0.25 17 15 11
0.20 28 20 15
0.15 45 30 25
0.10 89 50 48
0.05 201 170 147
d = 5 0.30 17 10 10
0.25 26 20 16
0.20 38 25 22
0.15 52 40 39
0.10 112 70 68
0.05 255 210 209
Table 2: Number of points necessary to reach the target star-discrepancies for Sobol’, subset selection and Scrambled Halton subsequences in dimensions 4 and 5
Since there are by far too many possible combinations of parameters in our method which cannot all be checked, we use the following algorithm to find $d$-dimensional sets with a particularly small star-discrepancy. Motivated by the bounds in Theorem 2, we take the $d$ smallest prime numbers. Then we define a numbers $n_{i}$ of shifts and $m_{i}$ of (randomly chosen) permutations to check for each dimension $i=1,\ldots,d$. Since the number of possible permutations grows faster than exponentially with the dimension, it is reasonable to increase the numbers $m_{i}$ with the $p_{i}$. Due to run time restrictions, $n_{i}$, i.e. the maximum value for the shift, is therefore defined decreasingly. To further decrease the number of possible permutations, we apply the heuristic from [BW79] and restrict ourselves to permutations $\pi_{i}$ such that $\pi_{i}(0)=0$ for all $i$.
For fixed $N\in\mathbb{N}$, our algorithm starts with dimension one, where $m_{1}=1$, because there is only one permutation for $p_{1}=2$, and search for the shift factor $1\leq a_{1}\leq n_{1}$ which minimizes the one-dimensional star-discrepancy. Then $a_{1}$ is fixed and we search for the optimal shift in the range $1\leq a_{2}\leq n_{2}$ (and permutation) such that the two-dimensional star-discrepancy is minimized. As $p_{2}=3$, there are only $2$ permutations with $\pi_{2}(0)=0$ which both can be checked. Next $p_{3}=5$ and the combination of shift $a_{3}$ and permutation $\pi_{3}$ which minimizes the three-dimensional star-discrepancy (given $a_{1},a_{2},\pi_{1},\pi_{2}$) is searched for. Then we fix $a_{3}$ and $\pi_{3}$, proceed with dimension $4$ and so on.
Table 2 compares the minimal $N$ required to find a set having star-discrepancy below a given threshold when using our algorithm (scrambled Halton), Sobol numbers and the subset selection algorithm. The numbers for the latter two methods are taken from Table 1 in [CDP23]. In dimension $d=4$, the scrambled Halton subsequences require a significantly lower number of points to achieve the listed target star-discrepancies than the other two methods. In dimension $d=5$, the differences are much smaller but the scrambled Halton subsequences are still superior. Note that even small changes can be relevant in practice, where one simulation (based on the assumptions defined by the point in the set) can take several hours and thus produce high computational and economic costs.
Moreover, our algorithm does not only yield smaller values of $N$, but has two other benefits in comparison to the subset selection algorithm: first, we can explicitly write down the parameters for the best found scrambled Halton subsequences in a compact form, so that the reader can independently reproduce our results. We report them for the results in Table 2 in Table 5 and Table 6 of the Appendix 4. Of course, it is also possible to write down the parametrization, i.e. the random seed, of the subset selection algorithm. However, in our eyes, this would be a less self-contained representation of the sets. Second, Halton subsequences are sequences and therefore after optimization for a given $N$, they can be extended to an arbitrary large $N_{1}>0$ by just using the given parametrization. We have seen in Figure 3 for the case of dimension $1$ that these sequences typically perform very well if $N_{1}$ is not too far apart from $N$.
Another approach to find sets with a particularly small star-discrepancy for a given $N$ was introduced in [DR13]. The main components of the algorithm therein are based on evolutionary principles. The algorithm is called optimized Halton and was the first algorithm which could be adapted easily to optimize the inverse star-discrepancy. Therefore, it can serve as another good point of comparison. For completeness, we also add to it the results from [DGW10], which served as a benchmark in [DR13].
In Table 3, we see that the scrambled Halton subsequences approach yields promising values for the best found star-discrepancy given the dimension $d$ and the set size $N$. While the star-discrepancy of the scrambled Halton subsequences is the smallest for dimensions $d=5$ and $d=7$, it is slightly worse in the case $d=9$. Since the results are not deterministic but depend on the random choices of the permutations, the two approaches here and in [DR13] should in our opinion therefore be regarded as equally competitive. The results from [DGW10] on the other hand yield much larger values for the star-discrepancy of the best found set.
$d$ $N$ Results [DGW10] Results [DR13] Scrambled Halton
$5$ $95$ $\approx 0.11$ $0.08445$ $0.083796$
$7$ $65$ $0.150$ $0.1361$ $0.1346$
$7$ $145$ $0.098$ $0.08640$ $0.08573$
$9$ $85$ $0.170$ $0.1435$ $0.14515$
Table 3: Exact star-discrepancy results for the sequences from [DR13] and [DGW10] and scrambled Halton sequences.
Two Dimensions.
Two dimensions may be regarded as an intermediate case when using our approach: on the one hand, it can be treated as a two-dimensional minimization task for the star-discrepancy of Halton sequences. On the other hand, it may also be seen as one-dimensional minimization task for the (two-dimensional) scrambled Hammersley point set defined by $y_{n}=(n/N,x_{n})\in[0,1)^{2}$ for $n=1,\ldots,N-1$, where $(x_{n})$ is a one-dimensional scrambled van der Corput subsequence. We will discuss the theoretical properties of Hammersley point sets in more detail in Section 3.
Also the case of two dimensions has lately attracted particular attention: it has been treated with the machine learning approach fro, [RKB+24], which we have already mentioned. In [CDKP24], two-dimensional sets are constructed by considering two lists of coordinates and an assignment matrix, linking the list of coordinates to the points’ position. This way, the optimization of the star-discrepancy can also be formulated as a problem of finding an optimal permutation. The base coordinates considered therein are $y_{n}=(n/N,x_{n})\in[0,1)^{2}$ for $n=1,\ldots,N-1$, where $(x_{n})$ is a low-discrepancy sequence, as well. In particular the Kritzinger sequence and the Kronecker sequence for the golden ratio are taken for $(x_{n})$. In the latter case, the set is also called Fibonacci point set.
In our setting, it turned out that the Hammersley point set yields much smaller discrepancy values than two-dimensional Halton sequences. Therefore, the presentation in this article restricts itself to Hammersley point sets. To be precise, we considered the Hammersley point set for the $9$ smallest primes and the allowed number of shifts were $1,000,500$ and $100$ for the first three prime numbers and $50$ for all others. In order to give every prime the same chance to be the winner the numbers of random permutations are set such that the total number of tries is always $1,000$.
În Table 4, we compare our approach with the best results obtained by the two alternative methods. It turns out that the star-discrepancy of the scrambled Hammersley point set is smaller than for the Fibonacci point set. Nonetheless, it is outperformed by both alternative methods. Still it might be promising to use scrambled Hammersley point sets as base coordinates in the approach of [CDKP24]. Moreover, we see from Table 7 in the appendix, that the configuration $p=2$ and the shift $n=509$ is always the best found scrambled Hammersley point set for $n\geq 260$ and should be a particularly good candidate.
$N$ Fibonacci Scrambled Ham. Result [RKB+24] Best result from [CDKP24]
$20$ $0.105833$ $0.0800$ $0.00666$ $0.06219$
$50$ $0.049165$ $0.0395$ $0.2742$
$100$ $0.0026105$ $0.0200$ $0.0188$ $0.01492$
$180$ $0.015165$ $0.0131$ $0.0115$ $0.00901$
$220$ $0.012407$ $0.0100$ $0.0097$ $0.00737$
$260$ $0.010894$ $0.0098$ $0.0084$ $0.00640$
$350$ $0.008731$ $0.0075$ $0.004872$
$420$ $0.00728$ $0.0065$ $0.0058$ $0.00412$
$500$ $0.00611$ $0.0055$ $0.0052$
Table 4: Comparison of star-discrepancy for scrambled Hammersley point sets and results from [RKB+24] and [CDKP24] as reported in [CDKP24].
More related work.
Finally, we do not want to sweep under the rug that other choices for the permutations $\pi$ than those in [BW79] have been made by several authors, both in the one- and multidimensional. The excellent article [FL09] contains a comprehensive overview of many of these approaches, as well as numerical results that evaluate their performance in certain integration tasks. However, all of these methods are purely based on optimizing the permutation and do not include any shifts. Since all of these methods date prior to 2009, we decided to use more recent results as point of comparison in this article.
3 Theoretical background
In this section we present the proofs of our theoretical results. We start by showing Theorems 1 and 2. For this purpose we introduce concepts from $p$-adic discrepancy theory first, see e.g. [Mei68, Som22, Wei25] for details.
Bound for scrambled van der Corput subsequences.
Recall that the $p$-adic absolute value $|\cdot|_{p}$ is for $a=\frac{b}{c}$ with $b,c\in\mathbb{Z}\setminus\{0\}$ defined as $|a|_{p}:=p^{-m}$, where $m$ is the highest possible power with $a=p^{m}\frac{b^{\prime}}{c^{\prime}}$ and $(b^{\prime}c^{\prime},p)=1$. It turns out to be useful to regard the values $f(n)$ of a permutation polynomial $f:\mathbb{N}\to\mathbb{N}$ as a sequence in the $p$-adic integers $\mathbb{Z}_{p}=\{x\in\mathbb{Q}_{p}\,:\,|x|_{p}\leq 1\}$, where $\mathbb{Q}_{p}$ are the $p$-adic numbers, i.e. the completion of $\mathbb{Q}$ with respect to $|\cdot|_{p}$. For a sequence $(x_{n})\subset\mathbb{Z}_{p}$ and $N\in\mathbb{N}$, the $p$-adic discrepancy is defined as
$\delta_{N}^{(p)}(x_{n}):=\sup_{z\in\mathbb{Z}_{p},k\in\mathbb{N}}\left|\frac{\#\left(\textrm{Disc}_{p}(z,k)\cap\{x_{1},\ldots,x_{N}\}\right)}{N}-\frac{1}{p^{k}}\right|$
and the $p$-adic disc at center $z\in\mathbb{Z}_{p}$ of radius $p^{-k}$ is given by
$\textrm{Disc}_{p}(z,k):=\left\{x\in\mathbb{Z}_{p}\,:\,\left|x-z\right|_{p}\leq p^{-k}\right\}.$
The concept of $p$-adic discrepancy allows to quantify the degree of uniformity in $\mathbb{Z}_{p}$ in a similar way as real discrepancy theory does in $[0,1]$. The notion originally stems from [Cug62]. By the same argument as in the real setting, it is straightforward to see that $\tfrac{1}{N}\leq D_{N}(x_{n})\leq 1$ holds for all sequences $(x_{n})\subset\mathbb{Z}_{p}$ and all $N\in\mathbb{N}$. However, the lower bound is in fact also sharp for sequences. Unlike in the real case there is no $\log(N)$ term required in the numerator as was shown in [Bee69].
The proof of Theorem 1 now essentially follows the arguments in [Wei25] and relies on Hensel’s Lemma.
Lemma 5 (Hensel’s Lemma).
Let $f(x)\in\mathbb{Z}[X]$ and let $a\in\mathbb{Z}$ with $f(a)\equiv 0\bmod\ p^{k},$ but $f^{\prime}(a)\not\equiv\bmod\ p^{k}$. Then there exist $b\in\mathbb{Z}$ with $f(b)\equiv 0\bmod\ p^{k+1}$ and $a\equiv b\bmod\ p^{k}$.
Another key ingredient is the following theorem of Meijer.
Theorem 6 (Meijer, [Mei68]).
Let $(x_{n})\subset\mathbb{Z}_{p}$ be an arbitrary sequence and denote its $p$-adic discrepancy for $N\in\mathbb{N}$ by $\delta_{N}^{(p)}$. Let $D_{N}$ be the extreme discrepancy of the corresponding (real) sequence $(\varphi_{p}(x_{n}))\subset[0,1)$, where
$\varphi_{p}(z)=\sum_{i=0}^{\infty}\left\{a_{i}p^{-i-1}\right\}$
for $z\in\mathbb{Z}_{p}$ with $p$-adic expansion
$z=\sum_{i=0}^{\infty}a_{i}p^{i}$
with coefficients $0\leq a_{i}<p$. Then it holds that
$\delta_{N}^{(p)}<D_{N}<\delta_{N}^{(p)}\left(2+\frac{2(p-1)}{\log p}\log\left((\delta_{N}^{(p)})^{-1}\right)\right).$
Proof of Theorem 1.
If $f:\mathbb{N}\to\mathbb{N}$ is a permutation polynomial $\bmod\ p^{2}$, it is a permutation polynomial $\bmod\ p$ as well and there are no solutions to $f^{\prime}(x)\equiv 0\mod p$. Now let $y\in\mathbb{Z}/p^{k}\mathbb{Z}$ and denote by $\overline{f}$ the reduction of $f\mod p$ and by $\overline{y}$ the reduction of $y\mod p$. By assumption there is a simple root for the polynomial $\overline{f}(x)-\overline{y}\equiv 0\mod p$, so Hensel’s lemma inductively implies that there is a unique solution to $f(x)\equiv y\mod p^{k}$. In other words a solution of the equation $\overline{f}(x)-\overline{y}\equiv 0\mod p$ can be lifted to a solution $\bmod\ p^{k}$ for arbitrary $k\in\mathbb{N}$ by Hensel’s lemma.
As $f$ is a permutational polynomial mod $p^{k}$ for all $k$, every disk $\textrm{Disc}_{p}(z,k)$ contains $\lfloor Np^{-k}\rfloor$ or $\lfloor Np^{-k}\rfloor+1$ elements of the sequence. This property remains true after applying the permutation $\pi$ which only permutes the residue classes $\bmod\ p^{k}$. Hence, the $p$-adic discrepancy satisfies $\delta_{N}^{(p)}(f(n))\leq\frac{1}{N}$ for all $N\in\mathbb{N}$. Theorem 6 thus implies the claim after realizing that $\varphi_{b}(f(n))$ therein equals the definition of the scrambled van der Corput sequence in (1). ∎
Theorem 1 transfers to the multi-dimensional setting of Halton sequences. In this case, we can use a multi-dimensional generalization of Theorem 6. In order to do so we need to introduce the concept of $p$-adic discrepancy in several variables first. Let $P=(p_{1},\ldots,p_{d})$ be a vector of primes and define the ring
$\mathbb{Z}_{P}:=\mathbb{Z}_{p_{1}}\times\mathbb{Z}_{p_{2}}\times\ldots\times\mathbb{Z}_{p_{d}}.$
For a vector $K=(k_{1},\ldots,k_{d})$ of non-negative integers and $Z=(z_{1},\ldots,z_{d})\in\mathbb{Z}_{P}$ consider the neighborhoods
$\textrm{Disc}_{P}(Z,K):=\textrm{Disc}_{p_{1}}(z_{1},k_{1})\times\ldots\times\textrm{Disc}_{p_{d}}(z_{d},k_{d}).$
in $\mathbb{Z}_{P}$. The normalized Haar measure on $\mathbb{Z}_{P}$ is denoted by $\mu$ such that we obtain $\mu(\textrm{Disc}_{P}(Z,K))=\prod_{i=1}^{d}p_{i}^{-k_{i}}$. For a sequence $(x_{n})$ in $\mathbb{Z}_{P}$ its $p$-adic discrepancy is then defined by
$\delta_{N}^{(P)}(x_{n}):=\sup_{Z\in\mathbb{Z}_{P},k\in\mathbb{N}_{0}^{d}}\left|\frac{\#\left(\textrm{Disc}_{P}(Z,K)\cap\{x_{1},\ldots,x_{N}\}\right)}{N}-\mu\left(\textrm{Disc}_{P}(Z,K)\right)\right|.$
Finally, the mapping $\varphi_{P}:\mathbb{Z}_{p}\to\mathbb{R}^{d}$ is for $A=(a_{1},\ldots,a_{d})\in\mathbb{Z}_{P}$ given by
$\varphi_{P}(A)=(\varphi_{p_{1}}(a_{1}),\ldots,\varphi_{p_{d}}(a_{d})).$
Meijer proved in [Mei68] the following connection between the discrepancy in $\mathbb{Z}_{p}$ and $\mathbb{R}^{d}$.
Theorem 7 (Meijer, [Mei68]).
Let $P=(p_{1},\ldots,p_{d})$ be a vector of distinct primes and let $(x_{n})$ be a sequence in $\mathbb{Z}_{P}$. If $\delta_{N}^{P}$ denotes the $P$-adic discrepancy of $(x_{n})$ in $\mathbb{Z}_{P}$ and $D_{N}$ the extreme discrepancy of $\varphi_{P}(x_{n})$ in $\mathbb{R}^{d}$, then it holds that
$2^{-d}\delta_{N}^{(P)}\leq D_{N}\leq\delta_{N}^{(P)}\left(2\sum_{k=1}^{d}g_{d}+\log\left(\left(\delta_{N}^{(P)}\right)^{-1}\right)^{d}\prod_{k=1}^{d}\frac{2(p_{k}-1)}{\log(p_{k})}\right).$
In order to show Theorem 2 it thus suffices to prove that $\delta_{N}^{(P)}(x_{n})=\frac{1}{N}$ for scrambled Halton subsequences.
Proof of Theorem 2.
At first let us consider the sequence $n=(n,\ldots,n)\in\mathbb{Z}_{P}$ and an arbitrary $\textrm{Disc}_{P}(Z,K)$. By the Chinese remainder Theorem, there are either $\lfloor N/\prod_{i=1}^{d}p_{i}^{k_{i}}\rfloor$ or $\lfloor N/\prod_{i=1}^{d}p_{i}^{k_{i}}\rfloor+1$ elements of this sequence in $\textrm{Disc}_{P}(Z,K)$ (this property was also stated in [Mei68]).
Since $f_{i}(n)$ is a permutation polynomial $\bmod\ p_{i}^{k_{i}}$ for all $i=1,\ldots,d$ and all $k_{i}\in\mathbb{N}_{0}$ and $\textrm{Disc}_{P}(Z,K)$ was arbitrary, this property regarding the number of elements remains true for the sequence $f(n)=(f_{1}(n),\ldots,f_{d}(n))$. Finally, also the $\pi_{i}$ only permute the residue classes $\bmod\ p_{i}^{k}$ of the sequence elements which does neither affect the fact that the number of elements in $\textrm{Disc}_{P}(Z,K)$ is $\lfloor N/\prod_{i=1}^{d}p_{i}^{k_{i}}\rfloor$ or $\lfloor N/\prod_{i=1}^{d}p_{i}^{k_{i}}\rfloor+1$. Thus, the $p$-adic discrepancy is $\delta_{N}^{(P)}\leq\frac{1}{N}$.
To see that the $p$-adic discrepancy is not equal to 00 choose an arbitrary point $Z=(z_{1},\ldots,z_{d})\in\mathbb{Z}_{P}$ which is an element of the sequence. Then take $k_{1}$ large enough such that $\textrm{Disc}_{p_{1}}(z_{1},k_{1})$ does not contain any other element of the sequence. Thus $\delta_{N}^{(P)}\geq\frac{1}{N}$ and the claim follows. ∎
Now, we will no longer apply the $p$adic discrepancy theory, but proceed with the proof of Theorem 4. For small dimensions we will use a standard construction based on Halton sequences to ensure that the bound $D_{N}^{*}(x_{n})\leq 2.4631832\sqrt{\frac{d}{N}}$ is satisfied. These are the so-called Hammersley point sets which lift Halton sequences in dimension $d$ to sets in dimension $d+1$. Furthermore, our proof relies on the concept of bracketing numbers which we introduce afterwards as well.
Lifting of sequences.
Let $(x_{n})$ be a sequence of points in $[0,1)^{d-1}$ with star-discrepancy $D_{N}^{*}(x_{n})$. Then it is possible to construct a point set in $[0,1)^{d}$ which has almost the same $d$-dimensional star-discrepancy. By defining $y_{n}=(n/N,x_{n})\in[0,1)^{d}$ for $n=1,\ldots,N-1$, and $P=\{y_{1},\ldots,y_{N-1}\}$ we obtain
$D_{N}^{*}(P)\leq\max_{1\leq M\leq N}\frac{M}{N}D_{M}^{*}(x_{n})+\frac{1}{N},$
see [Nie92], Lemma 3.7. For the Halton sequence, this specific point set is known as Hammersley point set and the bound becomes
$D_{N}^{*}(P)\leq c_{d-1}\frac{\log(N)^{d-1}}{N}+\frac{1}{N}.$ (3)
Vice versa, if a bound of the form $D_{N}^{*}(P)\leq c_{d}\log(N)^{d-1}/N$ holds for point sets $P$ in dimension $d$, then a bound of the form $D_{N}^{*}(x_{n})\leq c_{d-1}^{\prime}\log(N)^{d-1}/N$ is true for all sequences in $(x_{n})$ in $[0,1)^{d-1}$, see [Rot54].
Bracketing numbers.
The concept of bracketing numbers serves as an important tool for finding bounds on the star-discrepancy, compare e.g. [GPW21] and [DGS05]. Here we follow the definition given in [Gne08]: Let $A\subset[0,1]^{d}$ and $\delta\in(0,1]$. A finite set of points $\Gamma\subset[0,1]^{d}$ is called a $\delta$-cover of $A$, if for every $y\in A$, there exist $x,z\in\Gamma\cup\left\{0\right\}$ such that $x\leq y\leq z$ (to be read component-wise) and $\lambda_{d}([0,z])-\lambda_{d}([0,x])\leq\delta$, where $\lambda_{d}$ denotes the $d$-dimensional Lebesgue-measure. For $x\leq y$ the interval $[x,y]$ is defined by $[x,y]:=[x_{1},y_{1}]\times[x_{2},y_{2}]\times\ldots\times[x_{d},y_{d}]$ and similarly for half-open intervals. The bracketing number $N_{[\ ]}(A,\delta)$ of $A$ is the smallest number of closed axis-parallel boxes (or brackets) of the form $[x,y]$ with $x,y\in[0,1]^{d}$, satisfying $\lambda_{d}([0,y])-\lambda_{d}([0,x])\leq\delta$, whose union contains $A$. Here, we will use bracketing numbers for proving Theorem 4.
Bounds on the star-discrepancy.
The main ingredient for the proof is the following Theorem 9 which is to a great extent analogous to Theorem 3.5 in [GPW21]. The main difference is the restriction to the case $d\geq 5$. The remaining cases $d\leq 4$ in the proof of Theorem 4 can then be covered by Halton sequences and some additional ideas to reduce the required number of computer calculations.
If $N$ is big in comparison to $d\leq 4$, then the theoretical bounds for the Halton sequence/Hammersley point set as a low-discrepancy sequence/point set yields a better bound than the one in Theorem 4 anyhow. Moreover, it is possible to find sets with a sufficiently small star-discrepancy in the finitely many exceptions. In small dimensions, we will explain and use a theoretically justifiable trick which helps to speed up the necessary computer calculations, see Remark 13.
The sources for the improvements in comparison to [GPW21] are threefold: first, we use Halton sequences to address the case $d\leq 4$ while only the case $d=1$ had been addressed separately before. Second, we apply the involved inequalities more carefully than in [GPW21] and by that obtain sharper bounds. Last but not least another key ingredient for our proof is the following result from [Gne24] which improved on the previously known best bounds for $d$-dimensional bracketing numbers.
Theorem 8 (Gnewuch, [Gne24], Theorem 2.9).
The cardinality of the optimal $\varepsilon$-bracketing cover can be estimated as
$N_{[]}(d,\varepsilon)\leq\frac{d^{d}}{d!}\left(\frac{1}{\varepsilon}\right)^{d}$
for $d\geq 3$.
In order to derive Theorem 4 we proceed as in [GPW21] and prove the following result from which it follows easily for $d\geq 5$.
Theorem 9.
Let $d,N\in\mathbb{N}$ with $d\geq 5$. Let $X=(X_{n})$ be a sequence of uniformly distributed, independent random variables on $[0,1]^{d}$. Then for every $c\geq 2.4631832$
$D_{N}^{*}(X)\leq c\sqrt{\frac{d}{N}}$
holds with probability at least $1-e^{-(1.6728349c^{2}-10.1495427)\cdot d}$ implying that for every $q\in(0,1)$
$D_{N}^{*}(X)\leq 0.7731673\sqrt{10.1495427+\frac{\log\left((1-q)^{-1}\right)}{d}}\cdot\sqrt{\frac{d}{N}}$
holds with probability at least $q$.
Proof.
Let $\mu\in\mathbb{N},\mu\geq 2$ be arbitrary and choose a $2^{-\mu}$-cover $\Gamma_{\mu}$ of minimum size. Applying Theorem 8 and the Stirling formula implies
$|\Gamma_{\mu}|\leq\sqrt{\frac{2}{\pi d}}e^{d}2^{\mu d}.$
As we want to avoid pure repetition, we will mention and use an intermediate result in the proof of Theorem 3.5 in [GPW21]. To do so, we need to introduce three auxiliary variables. First we set
$c_{\mu}:=\frac{1}{1-\sqrt{\frac{\mu+1}{2\mu}}}$
and for $\tau_{\mu}\geq 0$ define
$c_{1}:=\sqrt{4\tau_{\mu}\left(1+\frac{1}{3c_{\mu}}\right)}.$
Now we use the following lemma, which is essentially a combination of Lemma 3.4 and the beginning of the proof of Theorem 3.5. in [GPW21].
Lemma 10 ([GPW21]).
Let $X=(X_{n})$ be a sequence of uniformly distributed and independent random variables in $[0,1]^{d}$. Then for $c_{0}\geq 0$, the inequality
$D_{N}^{*}(X)\leq c_{0}\left(1+c_{1}c_{\mu}\sqrt{\frac{\mu}{2^{\mu}}}\right)\sqrt{\frac{d}{N}}$
is satisfied with probability at least
$1-\sqrt{\frac{2}{\pi d}}e^{-(2c_{0}^{2}-\mu+\sigma)d}\left(1+\frac{e^{-((\mu-\sigma)(\mu\tau_{\mu}-1)+(1-\log(2))\mu-\zeta-\sigma)d}}{1-e^{(-((\mu-\sigma)\tau_{\mu}-\log(2)))d}}\right),$ (4)
where
$\sigma=\mu-d^{-1}\log\left(\frac{2|\Gamma_{\mu}|}{\sqrt{\frac{2}{\pi d}}}\right)$
and
$\zeta=1+\log(2)+\frac{\log(2)}{d}$
In our situation the two variables from the lemma are $\sigma=\mu-\log(2^{\mu})-1-\frac{\log(2)}{d}$ and $\zeta=1+\log(2)+\frac{\log(2)}{d}$. In order to be able to ignore the exponential term in front of the bracket in (4), we set $c_{0}=\sqrt{(\mu-\sigma)/2}$. We now want to make sure that the expression in the bracket of (4) is $\leq\sqrt{\frac{\pi d}{2}}$. Note that $(\mu-\sigma)$ may be replaced by $2c_{0}^{2}$, which simplifies the calculation. For $\mu=13$ and $\tau_{\mu}=\log(2)/(\mu-\sigma)+0.02120108$ the desired inequality is satisfied and $c$ is equal to the claimed value. ∎
This allows us to complete the proof of Theorem 4.
Proof of Theorem 4.
For $d=1$, the set $P:=\{1/2N,3/2N,\ldots,(2N-1)/2N\}$ satisfies $D_{N}^{*}(P)=\frac{1}{2N}$ which is stronger than the claim. In the case $d=2$, the Hammersley point set in base $2$ satisfies
$D_{N}^{*}(\varphi_{2}(n))\leq D_{N}(\varphi_{2}(n))\leq\frac{7}{2N}+\frac{1}{2\log(2)}\frac{\log(N)}{N}$
which is smaller than $2.463\sqrt{\frac{2}{N}}$ for $N>1$. In dimension $d=3$, the Hammersley point set in base $2$ and $3$ satisfies
$D_{N}^{*}(\varphi_{2,3}(n))\leq\frac{3}{N}+\frac{1}{N}\left(\frac{1}{2\log(2)}\log(N)+\frac{3}{2}\right)\left(\frac{1}{\log(3)}\log(N)+2\right)$
which is smaller that $2.463\sqrt{\frac{3}{N}}$ for $N>28$. However, for $N\leq 28$, the actual star-discrepancy of the Hammersley point set can be calculated with the help of a computer (or also by hand if desired) to see that the claimed inequality is actually true for all $N\in\mathbb{N}$.
For the case $d=4$ the mentioned theoretical bounds on the star-discrepancy of the Halton sequences in Theorem 2 only guarantees $D_{N}^{*}(P)\leq 2.463\sqrt{\frac{4}{N}}$ for $N>3\cdot 10^{5}$, even when ignoring the linear term. This means that it is not feasible to calculate the exact star-discrepancies for the missing $N$, even with a very fast computer. However, there are improved bounds on the star-discrepancy of Halton sequences according to Theorem 2.1 from [Ata04], see also [FKP15].
Theorem 11 (Atanassov, [Ata04]).
Let $(H^{b}_{n})$ be the Halton sequence in the base $b=(b_{1},\ldots,b_{d})$. Then $D_{N}^{*}(H^{b}_{n})$ is bounded from above by
$\left(\frac{1}{d!}\prod_{j=1}^{d}\left(\frac{\lfloor b_{j}/2\rfloor\log(N)}{\log(b_{j})}+s\right)+\sum_{k=0}^{d-1}\frac{b_{k+1}}{k!}\prod_{j=1}^{k}\left(\frac{\lfloor b_{j}/2\rfloor\log(N)}{\log(b_{j})}+k\right)\right)\frac{1}{N}.$
Applying Theorem 11 and (3) implies that the Hammersley point set satisfies $D_{N}^{*}(P)\leq 2.463\sqrt{\frac{4}{N}}$ for $N>11,759$. The finitely many (relatively few) excluded point sets can now be checked on a computer to satisfy the desired inequality. This completes the proof. ∎
Remark 12.
Note that our method could be extended to higher dimensions. For the case $d=5$ the theoretical bounds on the star-discrepancy of the Halton sequences from Theorem 11 guarantee $D_{N}^{*}(P)\leq 2.463\sqrt{\frac{5}{N}}$ for $N>7,800$. Again, these finitely many exceptions could be checked on a computer. However, the running time of the DEM algorithm is $O(N^{d/2+1})$, see [CVN+23]. Therefore, these finitely many exception become less and less feasible to check as the dimension increases.
Remark 13.
The following triangle inequality for the star-discrepancy accelerates the computer computation by eliminating the need to calculate star-discrepancies for many values of $N$.: Suppose that $D_{N_{0}}^{*}(X)$ is known and that we want to calculate $D_{N_{1}}^{*}(X)$ for $N_{1}=\alpha N_{0}$ with $\alpha>1$. Then according to [KN74], Theorem 2.6 on p. 115, it holds that
$D_{N_{1}}^{*}(X)\leq\frac{1}{\alpha}D_{N_{0}}^{*}(X)+\frac{\alpha-1}{\alpha}.$
Thus the star-discrepancy can increase by at most $\frac{\alpha-1}{\alpha}$. If $b=2.463\sqrt{\frac{d}{N}}-D_{N_{0}}^{*}(X)$, then we may choose $\alpha=1/(1-b)$. For instance for $N_{0}=5,000$ and $d=4$ we have $D_{N_{0}}^{*}(X)=0.0045$ and $0.06932.463\sqrt{\frac{d}{N}}=$ 0.0693 for the Hammersley point set in base $b=(2,3,5)$. Hence we could directly jump to $N_{1}=5,363$ elements and calculate their discrepancy if we dealt with a sequence. The gain is that it is not necessary to calculate any discrepancy in the range $3625,001\leq N\leq 5,$ 362. This trick decreases the necessary computational effort by a three-digit factor.
For the Hammersley point set, we need however to account for the fact that it is not a sequence: we know that the star-discrepancy of the underlying Halton sequence is at most $D_{N_{0}}^{*}(X)$ as well and the presented argument works for the Halton sequence when moving from $N_{0}$ to $N_{1}$ For the Hammersley point set with $N_{1}$ points, the star-discrepancy might be bigger by $1/N_{1}<1/N_{0}$ than the one of the Halton sequence according to (3). Hence we may only choose $\alpha=(1-1/N_{0})/(1-b)$. In our numerical example this means that we may only jump to $,35$ 353, which is however an acceptable difference.
It was conjectured in [NW21] that $n=10d$ points suffice to reach $D_{N}^{*}=0.25$. Our result constitutes partial progress towards this conjecture by a simple application of Theorem 4.
Corollary 14.
For every $d\in\mathbb{N}$ there exists a point set $S_{1}$ with $N=98d$ elements such that $D_{N}^{*}(S_{1})\leq 0.25$. Moreover, there exists a point set $S_{2}$ with $N=10d$ elements with $D_{N}^{*}(S_{2})\leq 0.7789269$.
Proof.
According to Theorem 4 for every $N,d\in\mathbb{N}$, there exists a point set $P$ such that
$D_{N}^{*}(P)\leq 2.4631832\sqrt{\frac{d}{N}}.$
Inserting $N=98d$ and $N=10d$ yields the claim. ∎
Remark 15.
With the value $c=2.4968$ from [GPW21], we would only obtain a set size $N=100d$ for $S_{1}$ and $D_{N}^{*}(S_{2})\leq 0.78956$.
4 Appendix
In the appendix, we report the parametrization of the algorithms as well as the scrambled Halton subsequences which yield the best found values for the inverse star-discrepancies (as presented in Tables 2 and 3). In Table 5 and 6 we used the parametrization $n_{\textrm{shifts}}=(100,100,40,40,40)$ and $m_{\pi}=(1,5,20,80,80)$. The data therein needs to be read as follows: at first we list the prime base $b$ for the Halton sequence (these are typically the first $k$ prime numbers). Afterwards we note the applied shifts $x_{i}\mapsto n_{i}\cdot x_{i}$ as vector $n$. Finally, we write down the permutation $\pi_{i}$ for each base element $b_{i}$ in the same ordering as the prime numbers. Thereby $\pi_{i,j}$, the $j$-th entry of $\pi_{i}$, is to be read as $\pi_{i}(j-1)=\pi_{i,j}$.
Target $D_{N}$ $N$ Data
0.30 8 $b=(2,3,5,7):n=(9,28,29,3);$
$\pi:(0,1);(0,2,1);(0,2,3,4,1);(0,5,3,1,2,4,6)$
0.25 11 $b=(2,3,5,7):n=(17,28,9,4);$
$\pi:(0,1);(0,2,1);(0,3,2,1,4);(0,1,2,6,3,5,4)$
0.20 15 $b=(2,3,5,7):n=(1,28,17,2);$
$\pi:(0,1);(0,2,1);(0,1,2,4,3);(0,3,4,1,2,5,6)$
0.15 25 $b=(2,3,5,7):n=(41,97,38,16);$
$\pi:(0,1);(0,1,2);(0,4,1,2,3);(0,3,5,4,1,6,2)$
0.10 48 $b=(2,3,5,7):n=(85,83,4,26);$
$\pi:(0,1);(0,1,2);(0,2,4,3,1);(0,4,3,2,6,1,5)$
0.05 147 $b=(2,3,5,7):n=(17,49,21,11);$
$\pi:(0,1);(0,2,1);(0,3,1,2,4);(0,5,2,6,4,3,1)$
Table 5: Parameters for the best found Halton subsequences in dimension $d=4$
Target $D_{N}$ $N$ Data
0.30 10 $b=(2,3,5,7,11):n=(29,73,21,20,9);$
$\pi:(0,1);(0,1,2);(0,1,4,2,3);(0,2,4,1,3,6,5);$
$(0,5,9,1,3,6,10,8,7,2,4)$
0.25 16 $b=(2,3,5,7,11):n=(17,28,21,3,6);$
$\pi:(0,1);(0,2,1);(0,1,3,2,4);(0,2,1,3,6,5,4);$
$(0,10,7,5,3,8,1,6,4,9,2)$
0.20 22 $b=(2,3,5,7,11):n=(93,85,21,27,12);$
$\pi:(0,1);(0,1,2);(0,2,3,1,4);(0,2,6,5,1,3,4);$
$(0,6,9,3,8,5,4,10,1,7,2)$
0.15 39 $b=(2,3,5,7,11):n=(33,8,1,20,18);$
$\pi:(0,1);(0,2,1);(0,2,4,1,3);(0,2,4,5,6,1,3);$
$(0,2,1,10,5,7,9,4,8,3,6)$
0.10 68 $b=(2,3,5,7,11):n=(5,82,34,2,5);$
$\pi:(0,1);(0,1,2);(0,2,4,3,1);(0,1,3,2,6,4,5);$
$(0,7,6,8,4,3,9,1,5,2,10)$
0.05 209 $b=(2,3,5,7,11):n=(3,8,2,11,38);$
$\pi:(0,1);(0,1,2);(0,1,3,4,2);(0,1,2,3,6,5,4);$
$(0,2,10,7,1,4,5,8,9,6,3)$
Table 6: Parameters for the best found Halton subsequences in dimension $d=5$
$N$ Data
$20$ $p=5,n=12,\pi=(0,1,2,3,4)$
$50$ $p=2,n=765,\pi=(0,1)$
$100$ $p=5,n=1,\pi=(0,3,4,1,2)$
$180$ $p=2,n=253,\pi=(0,1)$
$220$ $p=23,n=26$
$\pi=(0,12,9,6,1,10,15,14,7,20,18,3,2,16,5,4,19,13,8,21,17,11,22)$
$260$ $p=2,n=509,\pi=(0,1)$
$350$ $p=2,n=509,\pi=(0,1)$
$420$ $p=2,n=509,\pi=(0,1)$
$500$ $p=2,n=509,\pi=(0,1)$
Table 7: Parameters for the best found scrambled Hammersley point sets in dimension $d=2$
Regarding the parametrization of the algorithm for obtaining the results in Table 3, we used the same one as for the Tables 5 and 6 in the dimensions $4$ and $5$. Admittedly, finding the optimal parametrization found in Table 8, required a bit more fiddling for dimension $d=7$ because whenever the prime number $p=17$ was used, the star-discrepancy became larger than in [DGW10]. Besides this peculiarity, the parametrization was $n_{\textrm{shift}}=(100,100,40,40,40,40,40)$ and $m_{\pi}=(1,5,20,80,80,60,60)$. In dimension $d=9$ we used the parametrization $n_{\textrm{shift}}=(100,100,40,40,40,40,40,20,20)$ and $m_{\pi}=(1,5,20,80,80,60,60,60,60)$. Finally, we present the optimal parametrization for the scrambled Hammersley point sets in dimension $d=2$ in Table 7.
$d$ $N$ Data
$5$ $95$ $b=(2,3,5,7,11):n=(65,82,4,15,26);$
$\pi:(0,1);(0,1,2);(0,1,4,3,2);(0,1,6,3,2,4,5);$
$(0,2,7,1,5,9,6,4,3,8,10)$
$7$ $65$ $b=(2,3,5,7,11,13,23):n=(65,82,37,34,15,30,1);$
$\pi:(0,1);(0,1,2);(0,4,1,2,3);(0,4,2,3,5,6,1);$
$(0,6,3,9,2,7,10,1,5,8,4);$
$(0,2,7,10,3,9,12,11,6,1,4,5,8);$
(0,17,8,18,6,9,22,2,7,20,3,1,11,4,21,(0,17,8,18,6,9,22,2,7,20,3,1,11,4,21$\ldots$( 0 , 17 , 8 , 18 , 6 , 9 , 22 , 2 , 7 , 20 , 3 , 1 , 11 , 4 , 21 , …
10,5,14,19,15,16,12,13,1$\ldots$ 10,5,14,19,15,16,12,13,1)… 10 , 5 , 14 , 19 , 15 , 16 , 12 , 13 , 1 )
$7$ $145$ $b=(2,3,5,7,11,13,19):n=(49,89,4,31,9,16,24);$
$\pi:(0,1);(0,2,1);(0,3,2,1,4);(0,5,2,4,1,6,3);$
$(0,8,3,10,4,5,9,1,2,6,7);$
$(0,10,11,4,3,8,9,2,6,1,7,12,5);$
$(0,16,18,13,9,5,12,2,3,6,17,11,10,8,1,15,4,14,7)$
$9$ $85$ $b=(2,3,5,7,11,13,17,19,23):n=(3,14,39,38,5,6,10,17,6);$
$\pi:(0,1);(0,1,2);(0,2,3,4,1);(0,5,4,6,1,2,3);$
$(0,9,5,8,6,1,4,7,3,2,10);$
$(0,11,4,6,3,5,9,10,1,8,2,7,12);$
$(0,13,16,6,5,10,3,15,1,14,8,7,2,9,4,11,12);$
$(0,5,10,18,4,11,2,12,17,14,3,8,7,9,6,15,16,13,1);$
(0,9,3,12,17,15,22,2,16,5,1,18,10,(0,9,3,12,17,15,22,2,16,5,1,18,10$\ldots$( 0 , 9 , 3 , 12 , 17 , 15 , 22 , 2 , 16 , 5 , 1 , 18 , 10 , …
19,11,4,14,6,21,7,13,20,8$\ldots$ 19,11,4,14,6,21,7,13,20,8)… 19 , 11 , 4 , 14 , 6 , 21 , 7 , 13 , 20 , 8 )
Table 8: Parameters for the best found Halton subsequences in comparison to the algorithm from [DR13].
Figure 4 pushes the results from Figure 2 to $N=5,000$ and it becomes visible that $\varphi_{5,\pi,a}$ still has about the same quality as $\varphi_{5,\pi}$ also for $n>1,000$. Both sequences clearly have a smaller discrepancy than the standard van der Corput sequence $\varphi_{5}$.
Figure 4: Comparison of $\varphi_{5},\varphi_{5,\pi}$ and $\varphi_{5,\pi,1,000}$ up to $N=5,000$.